OpenStudy (howard-wolowitz):

For the polynomial function .............. find the zeros. Then determine the multiplicity at each zero and state whether the graph displays the behavior of a touch or a cross at each intercept.

OpenStudy (howard-wolowitz):

\[f(x)= -x ^{6} + 3x ^{4} + 4x ^{2}\]

OpenStudy (howard-wolowitz):

x = 0, touch; x = −2, cross; x = 2, cross x = 0, cross; x = −2, cross; x = 2, cross x = 0, cross; x = −2, touch; x = 2, touch x = 0, touch; x = −2, touch; x = 2, touch

OpenStudy (howard-wolowitz):

Hey!

OpenStudy (howard-wolowitz):

@Ilovecake

OpenStudy (ilovecake):

Sorry I don't know how to do this.....

OpenStudy (howard-wolowitz):

ok well thxs for trying

OpenStudy (anonymous):

ummmm............. I think it's #2

OpenStudy (howard-wolowitz):

Did you guesss?

OpenStudy (anonymous):

no of course not I think it's #2

OpenStudy (anonymous):

Then you guessed. It's 1.

OpenStudy (howard-wolowitz):

Thanks K

OpenStudy (howard-wolowitz):

dang what are yout typing?

OpenStudy (triciaal):

not the fastest way to do this the zero when f(x) = 0 f(x) = -x^6 + 3 x^4 + 4 x^2 = 0 -x^2(x^4 - 3 x^2 -4) let x^2 = a -a(a +1)(a - 4) x^2 = 0, -1, 2, -2 f'(x) = -6 x^5 + 12 x^3 + 8 x f"(x) = -30 x^ 4 + 36 + 8

OpenStudy (howard-wolowitz):

making it the first one

OpenStudy (howard-wolowitz):

dont you agree?

OpenStudy (triciaal):

last line should be *36 x^2

OpenStudy (triciaal):

crosses at -2, 2

OpenStudy (howard-wolowitz):

and x=0

OpenStudy (howard-wolowitz):

so which is it then?

OpenStudy (triciaal):

i think 2 crosses at 0 also

OpenStudy (triciaal):

https://www.desmos.com/calculator

OpenStudy (triciaal):

correction did the graph touches at 0 crosses at -2 and 2

OpenStudy (triciaal):

did you see the graph? x = 0, touch; x = −2, cross; x = 2, cross is the correct answer.

OpenStudy (howard-wolowitz):

yes and thank-you

OpenStudy (howard-wolowitz):

https://mathway.com/

OpenStudy (ilovecake):

Did you figure out how to do it?